Accounts Payable = 820,000 pesos

Unearned Revenue = 40% of Owner's Capital 

Accounts Receivable = 10% of total Liabilities

Owner's Withdrawal = 350,000 pesos

Owner's Capital = 400% of Owner's Withdrawal

Biological Assets = 540,000 pesos

Salaries Payable = 20% of Accounts Payable 

 

How much is the amount of Total Assets?


1
Expert's answer
2021-09-27T11:14:44-0400

Assets.

Accounts Receivable = 10% × 1544000=154400

Biological Assets =540,000 pesos

154400 + 540000= 694,400 pesos


Liabilities.

Accounts Payable =820,000 pesos

Salaries Payable = 20% × 820,000 =164,000

Unearned Revenue = 40% × 1400000 =560,000

Total Liabilities = 1544000


Owners' Equity

Owner's Capital = 400% ×350,000 =1400000

Owner's Withdrawal =350,000


Total Assets = Liabilities + Owner's Equity

Owner's Equity = Assets - liabilities

694,400 - 1544000 = (849,600)

Total Owner's Equity = (849,600) + 1400000+350,000 =900,400 pesos


Total Assets = Liabilities + Owner's Equity

= 1544000 + 900,400 = 2,444,400 pesos.